1.a.Solve the IVP 2 [:] = [3 =:] [:] []=[3] ņ

b.

Find e At where 2 5 A = -2 -4

c.

Solve the IVP (21-1 -3) M (O)-() x(0) I g(0)
Answer #1

Solution: x(0) la. + :11] 3 - 3 yo) Denote A = 3 det det (A-XI) - 2-1 -8-1 (2-2)(-8-2) +21 =-16-21 +81 +12 +21 = d²+67+5 7? ++ ést • 4 = e, et Truet +cest 3c, et + c est Excool 76 + 361 +₂ [sci teleb -P] Y(0) 76 t = 3 -34, + C₂ = 7 4C ==1 Tet_4est 24{(3+i) ]A=:] V₂ is free. Put ₂ = t :. v = (-3/2 - 1/2) = --3- Put t=2 7 eigenvector N] =:19] [:]-[:]+[:] [2 . to and D = -[3 sint + cost { (-3 cost +sint) - 1-3 sint-cost) {(2 cost)-(2 sint) -zsint (+) cost + ( + 월 ) sint rét sint+cost -zsint 2) sNow, we solue (1+35) = [:] · 3=] Its matrix is augmented -1 VS Vs .. Iz is free. Put S=t 6:51:] **: 1116 BIN M :. $. - $ =